Eq. wt. of Na2S2O3 in reaction:
2NaS2O3 + I2 ---> Na2S4O6 + 2NaI

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Hiii Pushpa
here 2 mole of NaS2O3 is given , 
so for one mole it will be – 2/2 = 1 
E quivalent mass = Molar mass / change in oxidtion no. 
So, = 158/1 = 158 
Hope this will help.